# Millinovae: a new class of transient supersoft X-ray sources without a nova eruption

### v1

@@ 5 Discussion and Conclusions
Overall, the optical (light curve shape, narrow emission lines, strong He ii emission) and X-ray properties of OGLE-mNOVA-11 match well those of ASASSN-16oh. We thus propose that they, and probably the remaining objects described in this study, form a homogeneous group of transient supersoft X-ray sources. We dub them ‘millinovae’ because their optical luminosities are roughly a thousand times fainter than those of ordinary classical novae. 111They should not be confused with ‘micronovae’, a much lower luminosity and much faster event seen in a number of Galactic magnetic cataclysmic variables (Scaringi et al., 2022).
@@
Their sky location (upper panels of Figure 4) and the fact that most of the selected stars occupy a relatively narrow region in the color–magnitude diagram (middle-left panel of Figure 4) indicate that millinovae are all in the Magellanic Clouds. We used the long-term astrometric time-series of OGLE observations to measure proper motions of selected objects (Appendix D), which are consistent (with one exception) with those of stars located in the Magellanic Clouds (Gaia Collaboration et al., 2018). The radial velocities of ASASSN-16oh (Maccarone et al., 2019) and OGLE-mNOVA-11 (Appendix B) are consistent with the systemic velocities of the SMC and LMC, respectively. However, we cannot rule out that some objects are located in the foreground Milky Way disk or in background galaxies, although we consider this very unlikely.
@@
The middle-left panel of Figure 4 shows that most objects in quiescence occupy the same region of the color–magnitude diagram, , , which indicates that the accretion disk dominates their quiescent luminosity. Such a range is consistent with the absolute magnitudes of dwarf novae in the Magellanic Clouds (Shara et al., 2003), indicating that they are essentially nova-like disks. In outburst (middle-right panel of Figure 4), objects move toward bluer colors () and higher luminosities (), as expected given the strong irradiation of the disk by the supersoft component.
@@
The mean absolute magnitudes of millinovae in quiescence are and . In outburst, the absolute magnitudes are and , and are about a thousand times fainter than classical novae (lower-right panel of Figure 4). Millinovae are also much brighter than most known dwarf novae (lower panels of Figure 4).
@@
The spectroscopic properties of ASASSN-16oh and OGLE-mNOVA-11, and the color evolution of OGLE-mNOVA-11 indicate that their optical outbursts likely have a disk origin. (If the optical outburst of OGLE-mNOVA-11 had a classical nova origin, the -band light curve would decline earlier than the -band; Hillman et al. 2019.) Because dwarf novae follow a well-known relation between the orbital period and absolute magnitude (both in quiescence and outburst; Patterson 2011), relatively high absolute magnitudes of millinovae indicate that their orbital periods should be on the order of a few days. Indeed, one of the analyzed objects (OGLE-mNOVA-08) shows clear eclipsing variability with a day period (Figure 5). Furthermore the rate of decline from outburst of dwarf novae, , is also known to be proportional to (the ‘Bailey relation’, Bailey 1975; see Figure 6), as expected given that longer periods imply larger disks. Most millinovae have decline rates in the range day mag-1, which corresponds to orbital periods from 3 to 15 day, assuming that the Bailey relation can be extrapolated to such long periods.
@@
The “spreading layer” (Kippenhahn & Thomas, 1978; Inogamov & Sunyaev, 1999; Piro & Bildsten, 2004; Maccarone et al., 2019) model can naturally produce the soft X-ray temperatures that have been seen in many dwarf nova outbursts, albeit in shorter period systems such as U Gem and SS Cyg (see e.g. Wheatley et al. 2003). However, the long orbital periods of millinovae imply the existence of large disks and hence likely large mass transfer rates (e.g., Echevarría 1994) capable of producing higher X-ray luminosities. Nevertheless, there are difficulties with this model that have been noted (e.g., Kato et al. 2020) in that there are high mass transfer rate recurrent novae in quiescence (e.g. U Sco, RS Oph) where such a soft component would have been expected, but has never been seen.
@@
The most obvious solution to explaining the properties of millinovae is then to find a way to trigger thermonuclear burning on the white dwarf surface without producing a nova flash. One such attempt to do this is the non-ejecting thermonuclear runaway event of Hillman et al. (2019), but the resulting light curves of their ASASSN-16oh model are completely different to the smooth rise and fall that is observed, not to mention the huge color variation they predict which is definitely excluded by the data.
@@
An alternative route to such a trigger has been sought by Kato et al. (2020) who build on the already known ongoing high mass-transfer rate in these systems (their brightness and variability properties are consistent with nova-likes in the Magellanic Clouds; Shara et al. 2003), thereby keeping the disk (and white dwarf) in a hot, active state. Furthermore, they require a massive white dwarf, close to 1.2 M⊙ in order to initiate steady burning without a strong wind or shell being ejected. While we do not yet know how many of these long symmetric outburst systems produce detectable supersoft X-rays, if they were all to do so then this would have important implications for the theory of non-ejecting supersoft X-ray sources, as it would constrain the parameters of systems capable of behaving in this way. Furthermore, it would provide a route for allowing the white dwarf to continue to grow in mass, and hence become a potential Type Ia supernova progenitor.
@@
We recognise, however, that an added complication to the Kato et al. (2020) models is that they produce close to Eddington-limited events, whereas some of our millinovae clearly do not reach such a level (Appendix E). Either there are severe geometrical constraints on the fluxes observed, or the nuclear burning does not cover the entire white dwarf surface. Nevertheless, we believe that this new group of millinovae, all likely representing long period, high mass transfer rate cataclysmic variables opens an important new route for study, with the added benefit of a well-constrained population in the Magellanic Clouds.
@@ Appendix A Light Curves
Figures 7 and 8 show the light curves of the remaining stars showing long-duration symmetrical outbursts, which we classified as ‘millinovae’.
@@ Appendix B SALT Observations of OGLE-mNOVA-11
Optical spectra of OGLE-mNOVA-11 were obtained with the Robert Stobie Spectrograph (Kobulnicky et al., 2003; Burgh et al., 2003) on the 10-m Southern African Large Telescope (SALT; Buckley et al. 2006) in long-slit mode. Two spectra were taken each night on 2023 November 22 and December 2, 3, 6, 8 and 10 with an exposure time of 1200 s for each spectrum. Observations were made with the PG0700 grating at an angle of with a slit width of , giving a resolving power of . The data were pre-reduced using the PySALT package (Crawford et al., 2010), which includes bias subtraction, gain and cross-talk corrections, and mosaicing. An Argon lamp was utilized for wavelength calibrations and the spectrophotometric standard star, EG 21, was used for flux calibration. The IRAF longslit package was used to perform the wavelength and flux calibrations, cosmic ray removal and spectra extractions. The spectra were stacked so as to obtain one average spectrum per night.
The software package, Fityk (https://fityk.nieto.pl/) (Wojdyr, 2010), was used for spectral analysis. The SALT spectra were normalized before the emission lines were modelled using the method of Shafter (1983) in order to determine the central wavelengths of the He ii (4686 Å), , and emission. This entailed fitting a Gaussian function through 95% of the line wings. In Fityk, the errors of the fitted parameters were obtained for a confidence level of (68.27%). The measured central wavelengths were then utilized to calculate the barycentric-corrected radial velocities.
Table 2 presents the barycentric radial velocities of He ii (4686 Å), H, and H. We also calculated the weighted mean radial velocity from He ii and H (the H line was detected at low significance, hence, the large uncertainties of its radial velocities). The measured values vary from to km s-1, with a mean of km s-1, which is very close to the established LMC line-of-sight velocity ( km s-1; van der Marel et al. 2002). The number of radial velocity measurements prevents us from determining the orbital period unequivocally, but it is likely to be shorter than 4 day. If we restrict the period to be longer than 1 day and assume a circular orbit, the best-fit periods are 2.42, 1.68, 2.08, and 1.89 day.
@@ C.1 Swift X-ray data
The Neil Gehrels Swift Observatory (Gehrels et al., 2004) performed five Target of Opportunity observations of OGLE-mNOVA-11 (Target ID 16408) between 2023 December 6 and December 30, collecting ks during each pointing, with the exception of December 17, when only ks was obtained. The X-ray Telescope (XRT; Burrows et al. 2005) was operated in Photon Counting mode, and the data were processed using HEASoft 6.32, together with the most up-to-date calibration files available. Given the faintness of the source, a small extraction region of 10 pixels () radius was used for the source, and a larger, 60 pixel () radius circle, offset from, but close to, the source was used to estimate the background contribution.
In observations where the source was detected at the 3-sigma level, the 0.3–10 keV count rate (corrected for all PSF losses) is given in Table 3. In the shorter exposure on December 17, the source was not detected, and a upper limit is given instead.
There is no significant change in the hardness ratio during these observations, though it should be noted that we are dealing with a limited number of source counts ( in total across all five observations, the vast majority below 1 keV). Therefore, all the data were combined to create a single spectrum. Cash statistics (Cash, 1979) were used to fit this spectrum with a combination of an optically-thick blackbody for the soft emission, and an optically-thin thermal plasma to account for the few harder energy photons, both absorbed by the same hydrogen column. While the optically-thin plasma can only be constrained to have a temperature of 4.2 keV ( K), the best-fitting blackbody temperature is 36 eV (418,000 K), absorbed by a column density of cm-2. This fit results in a C-stat value of 19.9 for 25 degrees of freedom.
If is fixed to the LMC value of cm-2 (HI4PI Collaboration et al., 2016), the blackbody temperature becomes slightly better constrained, with eV (temperature 607,000 K), for C-stat/dof = 21.2/26. This fit leads to an observed luminosity (in the energy range 0.3–10 keV, and at a distance of 49.59 kpc; Pietrzyński et al. 2019) of erg s-1, or erg s-1 unabsorbed.
@@ C.2 Swift UV data
We then estimated the brightness of the neighbor from the analysis of the UV–optical color–magnitude diagrams, assuming that the neighbor’s color is consistent with colors of other main-sequence stars of similar magnitude. We, therefore, created UV–optical color–magnitude diagrams by cross-matching the UVOT catalogs with the OGLE photometric map. We then calculated the mean color of main-sequence stars in the brightness range , where mag. We found:
(We also tested and and got similar results.) That enabled us to estimate the brightness of the neighbor in the UV filters:
@@ Appendix D OGLE Proper Motions
We calculated the proper motions of detected stars using the OGLE time-series astrometric data (OGLE-Uranus; A. Udalski et al. 2024, in preparation). In short, we measured each star’s position in individual OGLE images collected between 2010 and 2024 (in some cases 2010–2020 if later data were unavailable). We then cross-matched the positions of bright stars with those calculated using the Gaia Data Release 3 (DR3) data (Gaia Collaboration et al., 2016, 2023), taking into account their proper motion and heliocentric parallax. That allowed us to find the transformation between the positions in individual OGLE images to the equatorial coordinates in the Gaia reference frame (Lindegren et al., 2021). The proper motions in right ascension and declination for all detected stars are reported in Table 5. Because most of our targets are very faint (), the proper motion uncertainties are large (the median error bar is 2 mas yr-1). Nonetheless, in all cases (but one), the proper motions are consistent within the quoted errors with the Magellanic Clouds’ proper motions (Gaia Collaboration et al., 2018). The only exception is OGLE-mNOVA-15 with a total proper motion of mas yr-1, making it most likely to be a foreground object, and we therefore treat it as a candidate millinova.
@@
| OGLE-mNOVA-01 | ||||
| OGLE-mNOVA-02 | ||||
| OGLE-mNOVA-03 | ||||
| OGLE-mNOVA-04 | ||||
@@
| OGLE-mNOVA-08 | ||||
| OGLE-mNOVA-09 | ||||
| OGLE-mNOVA-10 | ||||
| OGLE-mNOVA-11 | ||||
| OGLE-mNOVA-12 | ||||
@@
| OGLE-mNOVA-14 | ||||
| OGLE-mNOVA-15 | ||||
| OGLE-mNOVA-16 | ||||
| OGLE-mNOVA-17 | ||||
| OGLE-mNOVA-18 | ||||
| OGLE-mNOVA-19 | ||||
| OGLE-mNOVA-20 | ||||
| OGLE-mNOVA-21 | ||||
@@
| OGLE-mNOVA-23 | ||||
| OGLE-mNOVA-24 | ||||
| OGLE-mNOVA-25 | ||||
| OGLE-mNOVA-26 | ||||
| OGLE-mNOVA-27 | ||||
| OGLE-mNOVA-28 | ||||
| OGLE-mNOVA-29 |
@@
The first catalog of X-ray sources detected by the eROSITA telescope array (Predehl et al., 2021) aboard the Spektrum-Roentgen-Gamma (SRG; Sunyaev et al. 2021) satellite was recently published by Merloni et al. (2024). The number of exposures, and so the depth of the survey, depends on sky position. Thanks to the SRG scanning law, objects located near the ecliptic poles are observed most frequently. Coincidentally, the LMC is located near the south ecliptic pole.
We cross-matched our list with the eROSITA catalog (Merloni et al., 2024), but none of the objects from Table 1 were detected by eROSITA. The catalog of Merloni et al. (2024) contains X-ray sources located in the western Galactic hemisphere detected during the first six months of eROSITA operations (from 2019 December 12 to 2020 June 11, that is, ). The overlap with OGLE observations is relatively small because the OGLE-IV operations were suspended on 2020 March 18 () due to the COVID-19 pandemic. Nonetheless, we found that four objects (OGLE-mNOVA-06, OGLE-mNOVA-10, OGLE-mNOVA-19, and OGLE-mNOVA-24) were in outburst during the first eROSITA All-Sky Survey. All four objects were relatively faint and reached a peak -band magnitude of during that time. Assuming that the X-ray flux scales proportionally with peak -band flux, and taking OGLE-mNOVA-11 as a benchmark, the maximal expected X-ray flux is on the order of erg s-1 cm-2 in the 0.3–10 keV range. The expected flux in the eROSITA soft and medium energy bands (0.2–2.3 keV) is about 6% larger, erg s-1 cm-2. However, each object must have been scanned by eROSITA multiple times and the time-averaged flux is likely to be smaller.
We queried the eROSITA Upper Limits service (Tubín-Arenas et al., 2024) to get upper flux limits in the 0.2–2.3 keV range for all four objects discussed above. These limits are calculated at a 99.87% () one-sided confidence level assuming a power-law spectrum with a photon index and a column density cm-2. We obtained limits of erg s-1 cm-2 at the positions of OGLE-mNOVA-(06,10,19,24), respectively. We followed a procedure outlined in Tubín-Arenas et al. (2024) to estimate upper limits for the best-fitting absorbed blackbody model. Using the appropriate eROSITA calibration files and XSPEC version 12.14.0 (Arnaud, 1996), we estimated the energy-to-count conversion factor of cm2 erg-1, which is only 3.4% smaller than that calculated using the fiducial model in Tubín-Arenas et al. (2024). Therefore, the upper limits for our best-fitting model are only 3.4% larger than default ones. In the 0.2–5 keV range, the corresponding limits are erg s-1 cm-2. We thus conclude that the expected time-averaged flux is close to or below the current eROSITA limits. It is therefore possible that the selected stars may be detected in future eROSITA data releases once more data are processed, or may be detected in data from individual scans.

### v2

@@ 5 Discussion and Conclusions
Overall, the optical (light-curve shape, narrow emission lines, strong He ii emission) and X-ray properties of OGLE-mNOVA-11 match well with those of ASASSN-16oh. We thus propose that they, and probably the remaining objects described in this study, form a homogeneous group of transient supersoft X-ray sources. We dub them “millinovae” because their optical luminosities are roughly a thousand times fainter than those of ordinary classical novae.111They should not be confused with “micronovae,” a much lower-luminosity and much faster event seen in a number of Galactic magnetic cataclysmic variables (Scaringi et al., 2022).
Using ASASSN-16oh and OGLE-mNOVA-11 as prototypes, millinovae should fulfill the following criteria:
1.
They exhibit symmetrical, triangle-shaped outbursts in the optical bands (, ).
2.
Outbursts last from a month to several months, i.e., substantially longer than the typical outbursts of short-orbital-period dwarf novae.
3.
The peak absolute magnitude in the optical bands is between that of classical novae and short-orbital-period dwarf novae.
4.
The optical spectra show narrow (FWHM smaller than a few hundred km s-1) emission lines of H and He ii (that is, no signatures of mass ejection are seen in the spectra).
5.
Soft X-ray emission (declining with decreasing optical luminosity) is seen during the outburst.
Only two systems (ASASSN-16oh and OGLE-mNOVA-11) meet all these five criteria. The remaining objects, given the lack of dedicated spectroscopic and X-ray observations during outbursts, remain millinova candidates.
@@
Their sky location (upper panels of Figure 4) and the fact that most of the selected stars occupy a relatively narrow region in the color–magnitude diagram (middle left panel of Figure 4) indicate that millinovae are all in the Magellanic Clouds. We used the long-term astrometric time series of OGLE observations to measure proper motions of selected objects (Appendix D), which are consistent (with one exception) with those of stars located in the Magellanic Clouds (Gaia Collaboration et al., 2018). The radial velocities of ASASSN-16oh (Maccarone et al., 2019) and OGLE-mNOVA-11 (Appendix B) are consistent with the systemic velocities of the SMC and LMC, respectively. However, we cannot rule out that some objects are located in the foreground Milky Way disk or in background galaxies although we consider this very unlikely.
@@
The middle left panel of Figure 4 shows that most objects in quiescence occupy the same region of the color–magnitude diagram, , , which indicates that the accretion disk dominates their quiescent luminosity. Such a range is consistent with the absolute magnitudes of dwarf novae in the Magellanic Clouds (Shara et al., 2003), indicating that they are essentially nova-like disks. In outburst (middle right panel of Figure 4), objects move toward bluer colors () and higher luminosities (), as expected given the strong irradiation of the disk by the supersoft component.
@@
The mean absolute magnitudes of millinovae in quiescence are and . In outburst, the absolute magnitudes are and and are about a thousand times fainter than classical novae (lower right panel of Figure 4). Millinovae are also much brighter than most known dwarf novae (lower panels of Figure 4).
@@
The spectroscopic properties of ASASSN-16oh and OGLE-mNOVA-11 and the color evolution of OGLE-mNOVA-11 indicate that their optical outbursts likely have a disk origin. Because dwarf novae follow a well-known relation between the orbital period and absolute magnitude (both in quiescence and outburst; Patterson 2011), relatively high absolute magnitudes of millinovae indicate that their orbital periods should be on the order of a few days. Indeed, one of the analyzed objects (OGLE-mNOVA-08) shows clear eclipsing variability with a day period (Figure 5). (We searched for possible periodic brightness variations in quiescence and during outbursts in all 29 objects but detected a statistically significant signal only in the case of OGLE-mNOVA-08.)
Furthermore, the rate of decline from the outburst of dwarf novae, , is also known to be proportional to (the “Bailey relation”; Bailey 1975; see Figure 6), as expected, given that longer periods imply larger disks. We find that most millinovae have decline rates in the range day mag-1, which corresponds to from 3 to 15 days, assuming that the Bailey relation can be extrapolated to such long periods.
@@
The “spreading layer” (Kippenhahn & Thomas, 1978; Inogamov & Sunyaev, 1999; Piro & Bildsten, 2004; Maccarone et al., 2019) model can naturally produce the soft X-ray temperatures that have been seen in many dwarf nova outbursts, albeit in shorter-period systems such as U Gem and SS Cyg (see, e.g., Wheatley et al. 2003). However, the long orbital periods of millinovae imply the existence of large disks and hence likely large mass-transfer rates (e.g., Echevarría 1994) capable of producing higher X-ray luminosities. Nevertheless, there are difficulties with this model that have been noted (e.g., Kato et al. 2020) in that there are high mass-transfer-rate recurrent novae in quiescence (e.g. U Sco, RS Oph) where such a soft component would have been expected but has never been seen.
@@
The most obvious solution to explaining the properties of millinovae is then to find a way to trigger thermonuclear burning on the white dwarf surface without producing a nova flash. One such attempt to do this is the nonejecting thermonuclear runaway event of Hillman et al. (2019), but the resulting light curves of their ASASSN-16oh model are completely different to the smooth rise and fall that are observed, not to mention the huge color variation they predict, which is definitely excluded by the data.
@@
An alternative route to such a trigger has been sought by Kato et al. (2020), who build on the already known, ongoing high mass-transfer rate in these systems (their brightness and variability properties are consistent with nova-like disks in the Magellanic Clouds; Shara et al. 2003), thereby keeping the disk (and white dwarf) in a hot, active state. Furthermore, they require a massive white dwarf, close to 1.2 in order to initiate steady burning without a strong wind or shell being ejected. While we do not yet know how many of these long symmetric outburst systems produce detectable supersoft X-rays, if they were all to do so, then this would have important implications for the theory of nonejecting supersoft X-ray sources, as it would constrain the parameters of systems capable of behaving in this way. Furthermore, it would provide a route for allowing the white dwarf to continue to grow in mass and hence become a potential Type Ia supernova progenitor.
@@
We recognize, however, that an added complication to the Kato et al. (2020) models is that they produce close to Eddington-limited events, whereas some of our millinovae clearly do not reach such a level in X-rays (Appendix E). There may be severe geometrical constraints on the fluxes observed, or the nuclear burning might not cover the entire white dwarf surface. Moreover, some of these models (see also Kato et al. 2022) predict the highest, close-to-Eddington flux to be in the far UV range (rather than soft X-rays), which can be tested with dedicated follow-up observations of future millinova outbursts. Nevertheless, we believe that this new group of millinovae, all likely representing long-period, high mass-transfer-rate cataclysmic variables, opens an important new route for study, with the added benefit of being a well-constrained population in the Magellanic Clouds.
@@ Appendix A Light Curves
Figures 7 and 8 show the light curves of the remaining stars exhibiting long-duration symmetrical outbursts, which we have classified as “millinovae.”
@@ Appendix B SALT Observations of OGLE-mNOVA-11
Optical spectra of OGLE-mNOVA-11 were obtained with the Robert Stobie Spectrograph (Kobulnicky et al., 2003; Burgh et al., 2003) mounted on SALT (Buckley et al. 2006) in long-slit mode. Two 1200 s exposure spectra were taken each night on 2023 November 22 and December 2, 3, 6, 8, and 10, using the PG0700 grating at an angle of with a slit width of , giving a resolving power of . The data were prereduced using the PySALT package (Crawford et al., 2010), which includes bias subtraction, gain and cross-talk corrections, and mosaicking. An Argon lamp was utilized for wavelength calibration, and the spectrophotometric standard star, EG 21, was used for flux calibration, both performed with IRAF. The spectra were stacked so as to obtain one average spectrum per night.
The software package, Fityk222https://fityk.nieto.pl/ (Wojdyr, 2010), was used for spectral analysis. The SALT spectra were normalized before the emission lines were modeled using the method of Shafter (1983) in order to determine the central wavelengths of the He ii (4686 Å), , and emission lines. This entailed fitting a Gaussian function through 95% of the line wings. In Fityk, the errors of the fitted parameters were obtained for a confidence level of (68.27%). The measured central wavelengths were then utilized to calculate the barycentric-corrected radial velocities.
Table 2 presents the barycentric radial velocities of He ii (4686 Å), H, and H. Because of the larger scatter in our H velocities (due to its lower-significance detections), we also calculated the weighted mean radial velocity from He ii and H, obtaining values that ranged from to km s-1, with a mean of km s-1, which is very close to the established LMC line-of-sight velocity ( km s-1; van der Marel et al. 2002). The number of radial velocity measurements prevents us from determining the orbital period unequivocally, but it is likely to be shorter than 4 days. If we restrict orbital periods to be longer than 1 day and assume a circular orbit, the best-fit periods are 2.42, 1.68, 2.08, and 1.89 days.
@@ C.1 Swift X-ray Data
The Neil Gehrels Swift Observatory (Gehrels et al., 2004) performed five Target of Opportunity observations of OGLE-mNOVA-11 (Target ID 16408) between 2023 December 6 and December 30, collecting ks during each pointing, with the exception of December 17, when only ks was obtained. The XRT (Burrows et al. 2005) was operated in Photon Counting mode, and the data were processed using HEASoft 6.32, together with the most up-to-date calibration files available. Given the faintness of the source, a small extraction region of 10 pixels () radius was used for the source, and a larger, 60 pixel () radius circle, offset from, but close to, the source was used to estimate the background contribution.
In observations where the source was detected at the level, the 0.3–10 keV count rate (corrected for all point-spread function losses) is given in Table 3. In the shorter exposure on December 17, the source was not detected, and a upper limit is given instead.
There is no significant change in the hardness ratio during these observations though it should be noted that we are dealing with a limited number of source counts ( in total across all five observations, the vast majority below 1 keV). Therefore, all the data were combined to create a single spectrum. Cash statistics (Cash, 1979) were used to fit this spectrum with a combination of an optically thick blackbody for the soft emission and an optically thin thermal plasma to account for the few higher-energy photons, both absorbed by the same hydrogen column. While the optically thin plasma can only be constrained to have a temperature of 4.2 keV ( K), the best-fitting blackbody temperature is 36 eV (418,000 K), absorbed by a column density of cm-2. This fit results in a C-stat value of 19.9 for 25 degrees of freedom.
If is fixed to the LMC value of cm-2 (HI4PI Collaboration et al., 2016), the blackbody temperature becomes slightly better constrained, with eV ( 607,000 K), for C-stat/degrees of freedom = 21.2/26. This fit leads to an LX (0.3–10 keV) of erg s-1, or erg s-1 unabsorbed (assuming a distance of 49.59 kpc; Pietrzyński et al. 2019).
@@ C.2 Swift UV data
We then estimated the brightness of the neighbor from the analysis of the UV–optical color–magnitude diagrams, assuming that the neighbor’s color is consistent with colors of other main-sequence stars of similar magnitude. We, therefore, created UV–optical color–magnitude diagrams by cross-matching the UVOT catalogs with the OGLE photometric map. We then calculated the mean color of main-sequence stars in the brightness range , where mag. We found
(We also tested and and obtained similar results.) That enabled us to estimate the brightness of the neighbor in the UV filters:
@@ Appendix D OGLE Proper Motions
We calculated the proper motions of detected stars using the OGLE time-series astrometric data (OGLE-Uranus; A. Udalski et al. 2024, in preparation). In short, we measured each star’s position in individual OGLE images collected between 2010 and 2024 (in some cases, 2010–2020 if later data were unavailable). We then cross-matched the positions of bright stars with those calculated using the Gaia Data Release 3 (DR3) data (Gaia Collaboration et al., 2016, 2023), taking into account their proper motion and heliocentric parallax. That allowed us to find the transformation between the positions in individual OGLE images to the equatorial coordinates in the Gaia reference frame (Lindegren et al., 2021). The proper motions in right ascension and declination for all detected stars are reported in Table 5. Because most of our targets are very faint (), the proper motion uncertainties are large (the median error bar is 2 mas yr-1). Nonetheless, in all cases (but one), the proper motions are consistent within the quoted errors with the Magellanic Clouds’ proper motions (Gaia Collaboration et al., 2018). The only exception is OGLE-mNOVA-15 with a total proper motion of mas yr-1, making it most likely to be a foreground object, and we therefore treat it as a candidate millinova.
@@
| OGLE-mNOVA-01 | … | … | ||
| OGLE-mNOVA-02 | … | … | ||
| OGLE-mNOVA-03 | … | … | ||
| OGLE-mNOVA-04 | … | … | ||
@@
| OGLE-mNOVA-08 | … | … | ||
| OGLE-mNOVA-09 | … | … | ||
| OGLE-mNOVA-10 | … | … | ||
| OGLE-mNOVA-11 | … | … | ||
| OGLE-mNOVA-12 | … | … | ||
@@
| OGLE-mNOVA-14 | … | … | ||
| OGLE-mNOVA-15 | … | … | ||
| OGLE-mNOVA-16 | … | … | ||
| OGLE-mNOVA-17 | … | … | ||
| OGLE-mNOVA-18 | … | … | ||
| OGLE-mNOVA-19 | … | … | ||
| OGLE-mNOVA-20 | … | … | ||
| OGLE-mNOVA-21 | … | … | ||
@@
| OGLE-mNOVA-23 | … | … | ||
| OGLE-mNOVA-24 | … | … | ||
| OGLE-mNOVA-25 | … | … | ||
| OGLE-mNOVA-26 | … | … | ||
| OGLE-mNOVA-27 | … | … | ||
| OGLE-mNOVA-28 | … | … | ||
| OGLE-mNOVA-29 | … | … |
@@
The first catalog of X-ray sources detected by the eROSITA telescope array (Predehl et al., 2021) on board the Spectrum-Roentgen-Gamma (SRG; Sunyaev et al. 2021) satellite was recently published by Merloni et al. (2024). The number of exposures, and so the depth of the survey, depends on sky position. Thanks to the SRG scanning law, objects located near the ecliptic poles are observed most frequently. Coincidentally, the LMC is located near the south ecliptic pole.
We cross-matched our list with the eROSITA catalog (Merloni et al., 2024), but none of the objects from Table 1 were detected by eROSITA. The catalog of Merloni et al. (2024) contains X-ray sources located in the western Galactic hemisphere detected during the first six months of eROSITA operations (from 2019 December 12 to 2020 June 11, that is, ). The overlap with OGLE observations is relatively small because the OGLE-IV operations were suspended on 2020 March 18 () due to the COVID-19 pandemic. Nonetheless, we found that four objects (OGLE-mNOVA-06, OGLE-mNOVA-10, OGLE-mNOVA-19, and OGLE-mNOVA-24) were in outburst during the first eROSITA All-Sky Survey. All four objects were relatively faint and reached a peak during that time. Assuming that the X-ray flux scales proportionally with peak -band flux, and taking OGLE-mNOVA-11 as a benchmark, the maximal expected X-ray flux is erg cm-2 s-1 in the 0.3–10 keV range. The expected flux in the eROSITA soft and medium energy bands (0.2–2.3 keV) is about 6% larger at erg cm-2 s-1. However, each object must have been scanned by eROSITA multiple times, and the time-averaged flux is likely to be smaller.
We queried the eROSITA upper limits service (Tubín-Arenas et al., 2024) to get upper flux limits in the 0.2–2.3 keV range for all four objects discussed above. These limits are calculated at a 99.87% () one-sided confidence level, assuming a power-law spectrum with a photon index and a column density cm-2. We obtained limits of erg cm-2 s-1 at the positions of OGLE-mNOVA-(06,10,19,24), respectively. We followed a procedure outlined in Tubín-Arenas et al. (2024) to estimate upper limits for the best-fitting absorbed blackbody model presented in Appendix C. Using the appropriate eROSITA calibration files and XSPEC version 12.14.0 (Arnaud, 1996), we estimated the energy-to-count conversion factor of cm2 erg-1, which is only 3.4% smaller than that calculated using the fiducial model in Tubín-Arenas et al. (2024). Therefore, the upper limits for our best-fitting model are only 3.4% larger than default ones. In the 0.2–5 keV range, the corresponding limits are erg cm-2 s-1. We thus conclude that the expected time-averaged flux is close to or below the current eROSITA limits. It is therefore possible that the selected stars may be detected in future eROSITA data releases once more data are processed or may be detected in data from individual scans.